OpenStudy (anonymous):

rationalize the numerator and simplify 3-√x+2 / 21-3x please show workings! really need help with these :(

hero (hero):

Rationalize the numerator????????????

OpenStudy (anonymous):

or denominator!\

hero (hero):

The denominator doesn't need rationalizing

OpenStudy (anonymous):

so then how would i do this?

hero (hero):

Are you sure that's the original question?

OpenStudy (anonymous):

yeah sadly!

hero (hero):

It's a stupid question

hero (hero):

It's not your fault though

OpenStudy (anonymous):

haha yeah :P the questions the teacher gives us are really difficult, i got about 20 more questions in which i need help with :(

hero (hero):

I don't get what rationalize the numerator means

OpenStudy (anonymous):

yeah same here!

hero (hero):

\[\frac{3-\sqrt{x+2}}{21-3x}\] Is that it?

OpenStudy (anonymous):

yeah!

hero (hero):

multiply top and bottom by \[3 + \sqrt{3x+2}\] and see what happens

OpenStudy (anonymous):

alright thanks

hero (hero):

I'm going to go watch another episode of Smallville. I'll be back later